OpenStudy (anonymous):

If you burn 42.7 g of hydrogen and produce 382 g of water. How much oxygen reacted?

OpenStudy (jhannybean):

\[\color{red}{\sf 2H_2~(g)~ + ~O_2~(g) ~\rightarrow 2H_2O~(l)}\]You can either use the 42.7 g \(\sf H_2\) or you can use the 382 g \(\sf H_2O\) to find how much \(\sf O_2\) was used due to The Law of Conservation of Mass in a chemical reaction. It will approximately give you the same result. \[\sf 382 ~g~ H_2O ~ -~42.7~g ~H_2 = 339~g ~O_2 \] Let's check this using Stoichiometry. \[\sf 42.7~g~ H_2 ~\times~ \frac{1~mol~H_2}{2.02~g~ H_2} ~\times~ \frac{1~mol~O_2}{2~mol~H2}~\times~ \frac{31.9~g~O_2}{1~mol~O_2} \approx 338 ~g~O_2\]\[\sf 382~g~H_2O~\times~\frac{1~mol~H_2O}{18.0~g~H_2O}~\times~\frac{1~mol~O_2}{2~mol~H_2O}~\times~\frac{31.9~g~O_2}{1~mol~O_2}~\approx 339~g~ O_2 \] So you could see that they pretty much give you the same result when you subtract the two amounts.

OpenStudy (jhannybean):

@Shalante